In a coin bank there are 250 dimes and quarters worth a total of $39.25. Find the number of dimes and quarters in the bank.
my quess of equation but not sure:

d = dimes ; q = quarter ; c = coins
d + q = 250c ; dc + qc = $39.25
Am I on the right track?

Answer by mananth(16946)

dimes -------d
quarters ----------q
d+q= 250
10d+25c=3925 ( convert to cents)
multiply first equation by -10
-10d-10q=-2500
add to second equation
15q=1425
/15
q= 95
balance is dimes
250-95=155